Insta Limited Statement of Comprehensive Income For the year ended June 30, 2017 Amount Particulars in Rupees) Sales 843,500 Opening Inventory Purchases Factory Expenses Closing Stock Cost of Goods Sold 55,000 465,000 143,000 (79,500) (583,500) Gross Profit 260,000 Office Expenses Selling Expenses Distribution Expenses Provision for Doubtful Debts Depreciation Interest of Debentures 50,000) (30,000) (20,000) (4,000) (16,000) (14,000) Other Income Gain on disposal of vehicle Interest on Investment 20,000 40,000 Net Profit 186,000 Required: Based on the information given, compute the following ratios a) Gross Profit Margin 300.82 b) Operating Ratio c) Net Profit Margin
Answer #1

Ans : Gross profit margin = Gross profit / Total Revenue (Sales)

= 260,000 / 843,500

= 30.82 %

Operating Expenses = Cost of goods sold +Office expense +selling expense +Distribution expense +Provision for doubtful debts +Depreciation

= 583,500 +50000 + 30000 + 20000 + 4000 +16000

= 703,500

Operating ratio = operating expense / Net sales

= 703,500 / 843,500

= 83.4 %

Net profit margin = Net profit / Net Sales

= $ 186,000 / $ 843,500

= 22.05 %